OpenStudy (anonymous):

Find the distance between these points. C(0, 4), T(-6, -3)

OpenStudy (diyadiya):

Do you know the Distance formula?

OpenStudy (anonymous):

x1+x2 over 2 and y1+y2 over 2?

OpenStudy (diyadiya):

Nope Distance between two points (x1,y1) & (x2,y2) is \[\Large D= \sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

oh....

OpenStudy (diyadiya):

So what do you get?

OpenStudy (anonymous):

So would it be, (0-6)2 + (4-3)2?

OpenStudy (diyadiya):

(0, 4 ) (-6 ,-3) \[\sqrt{(-6-0)^2 + (-3-4)^2}\]

OpenStudy (anonymous):

so would I multiply -6 twice? this is where I get lost

OpenStudy (diyadiya):

(-6)^2 = -6 x -6

OpenStudy (anonymous):

so it'd be -6 as the outcome? and 49

OpenStudy (diyadiya):

No what is -6 x -6 ?

OpenStudy (anonymous):

36. so it would be 36+49? I did -3-4=-7, -7x-7=49

OpenStudy (diyadiya):

Yes But DON'T Forget the square root :D

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (diyadiya):

No problem :D